# Controllable Emotion Generation with Emotion Vectors

### v1

@@ Abstract
In recent years, technologies based on large-scale language models (LLMs) have made remarkable progress in many fields, especially in customer service, content creation, and embodied intelligence, showing broad application potential. However, The LLM’s ability to express emotions with proper tone, timing, and in both direct and indirect forms is still insufficient but significant. Few works have studied on how to build the controlable emotional expression capability of LLMs. In this work, we propose a method for emotion expression output by LLMs, which is universal, highly flexible, and well controllable proved with the extensive experiments and verifications [^1] Codes and data will be available at https://github.com/xuanfengzu/EmotionVector. This method has broad application prospects in fields involving emotions output by LLMs, such as intelligent customer service, literary creation, and home companion robots. The extensive experiments on various LLMs with different model-scales and architectures prove the versatility and the effectiveness of the proposed method.
Controllable Emotion Generation with Emotion Vectors
Yurui Dong1thanks: Equal Contribution, Luozhijie Jin1, Yao Yang2, Bingjie Lu2, Jiaxi ${}^{\text{\textdagger}}$, Zhi Liu3thanks: Corresponding author 1Fudan Unversity, 2Zhejiang Lab, 3Nanjing University of Chinese Medicine yuruidong22@m.fudan.edu.cn, jiaxiyang@zhejianglab.com, zhiliu@njucm.edu.cn
@@ 1 Introduction
Figure 1: When asking questions to a LLM, almost all models will answer the user’s question "politely" as shown in the figure, but when we apply our emotion vector, the model will produce strong emotional expressions. The example in the figure uses the llama3.1-8B-Instruct model and applies the extracted anger vector. More detailed examples are shown in Table 1.
In the field related with emotion, most NLP works have long been concentrating on the analysis and interpretation of human emotions, primarily through sentiment analysisDemszky et al. (2020); Gera et al. (2022); Zhang et al. (2024). These researches have provided valuable insights into understanding human language by categorizing text as different emotionsKim and Vossen (2021); Song et al. (2022). However, these works have largely overlooked an equally important aspect: how the models themselves might express emotionsMao et al. (2022).
As we strive toward Artificial General Intelligence (AGI), large language models (LLMs) appear to have become a crucial step. Some researches reveal that LLMs tend to exhibit a degree of self-cognitionChen et al. (2024a); Wang et al. (2023). However, this self-awareness often proves to be uncontrollable and prone to generating harmfulAndriushchenko et al. (2024), unlawful, or toxic outputsHartvigsen et al. (2022). As a result, developers typically align and suppress this self-cognition through reinforcement learningWang et al. (2024) or promptingGehman et al. (2020) to mitigate such risks, ensuring the models remain safe and aligned with human values.
Emotion, as one of the key representations of human self-cognition, still plays a critic role in controlling models’ outputLi et al. (2023). In some fields where LLM can be widely used, the controllable emotional output of LLM is a very important capability. For example, customer service requires a controllable emotional mechanism to ensure service qualityJo and Seo (2024), to avoid mechanical and stiff responses that affect the users’ experience. and content creators sometimes need to create texts with specified emotions. In embodied intelligence, the emotional expression ability of companion robots is the key point of customer experience. In the field of mental health care, there is a growing need for emotionally expressive models capable of providing emotional supportGrandi et al. (2024); Zheng et al. (2023) to enhance mental health outcomes.
Based on these challenges and requirements, we consider investigating how LLMs generate emotions and how to control it to be a highly important endeavor. We claim that LLMs inherently possess the capability to express emotions; but this ability has been suppressed as a result of strong alignment with human values. If we want to revoke the ability of models to deliver emotions, some stimuli need to be adapted, such as instruct tuningLiu et al. (2024b). While instruct tuning models show promising results, they often lack flexibility and fail to generalize across diverse applications and model architecturesGhosh et al. (2024). Some approaches rely on predefined emotion categories or assume a fixed set of emotional expressions, making them less adaptable to real-world, dynamic scenariosLiu et al. (2024b).
@@
In this paper, we propose an elegant but effective method for the controllable emotional and affective expressions LLMs. Our approach offers a universal solution that allows fine-grained control over the emotional tone and sentiment of generated text, without compromising its fluency or coherence. Our method only needs to use the prompt method to extract the "Emotion Vector" used by the LLM to express basic emotions. By applying EV in LLM’s inference process, we can achieve controllable adjustment of the emotion of the text generated by LLM and generate any answer with the emotion we want. Additionally, by demonstrating its effectiveness on a range of LLM architectures, our approach overcomes the limitations of previous methods that are tied to specific models or training sets.
2 Related Work
Emotional Dialog Systems
In order to create an agent or dialog system simulating the way that human beings express themselves, many studies was trying to find a way to make an emotional dialog system as emotion is the basic representation of human beings. Zhou et al. (2018) and Song et al. (2019) proposed a way of Emotion Embedding to make the model "has" the emotion, where, models were forced to install a module to generate emotions.
Instruct tuning based emotional control
@@ 2 Related Work
In the domain of emotional control and generation, a significant body of work has focused on leveraging fine-tuning techniques for LLMs. Chen et al. (2023) and Chen et al. (2024b) explored fine-tuning approaches to cultivate empathetic behavior in LLMs, specifically for applications within psychological counseling and emotional support domains. Furthermore, Zheng et al. (2023) proposed a specialized dataset and demonstrated how fine-tuning the Llama model could be used to create emotionally intelligent chatbots designed for empathetic interaction. However, althrough instruct-tuning models have relatively good performance, they are often inflexible and struggle to adapt to a wide range of applications and model architecturesGhosh et al. (2024). Some methods depend on predefined emotion categories or assume a fixed set of emotional expressions, limiting their ability to adjust to real-world, dynamic situations Liu et al. (2024b).
In-context Vectors and Function Vectors
Liu et al. (2024a) proposes the concept of In-context vector(ICV). ICV is added during forward propagation and used as a condensed contextual prompt. However, ICV only focuses on the last token position during extraction and lacks global significance. Similarly, Todd et al. (2024) proposes the Function Vector(FV). The FV they extracted pays more attention to the output of the attention head with the best average indirect effect, and then replaces the attention head at the corresponding position in the forward propagation to achieve improved performance of the model on specific tasks. The process of observing and extracting FV is relatively complicated. At the same time, since FV focuses on the process of causal analysis, it is difficult to apply to tasks such as emotions that require high generalization.Ilharco et al. (2023) also proposes a similar concept of task vector, but they need to fine-tune the model when extracting task vector, which is also a bit cumbersome compared to our method of directly using prompt to extract.
@@ 3.1 Constructing Emotion Vectors
To capture the emotional factors and semantics for LLM, a specialized dataset is designed and constructed to elicit specific emotional responses, referred to as EmotionQuery. The dataset consists of 500 queries, with 100 queries generated for each of five emotional states derived from the basic emotion modelsEkman (1992): joy, anger, disgust, fear, and sadness to provoke the corresponding emotional reactions. The queries were generated by a GPT-4o-miniOpenAI (2024). A more detailed description of the dataset and query construction process can be found in the Appendix B.1.
@@
$\Delta O_{l}^{e_{k}}=\bar{O}_{l}^{\text{emotion}(e_{k})}-\bar{O}_{l}^{\text{neutral}},$
@@ 3.2 Steering Emotion Vectors
$EV^{\text{base}}=\frac{1}{K}\sum_{k=1}^{K}EV^{e_{k}},$
@@
Figure 2: The pipeline of how we steer the Emotion Vectors.
@@ 4 Experiments
To evaluate the effectiveness of our proposed emotion vectors (EVs), we designed experiments to assess three key aspects: (1) whether adding EVs successfully imbues the model’s outputs with emotional tone, and (2) whether the application of EVs affects the original semantics and fluency of the generated sentences. (3) whether applying a scalar factor to the EVs improves the emotional intensity or tone. Specifically, we constructed a new dataset, EmotionQuery+ (EQ+), which is described in detail in Appendix B.2. This dataset includes 50 queries for each of the five emotional states from the EmotionQuery dataset, along with an additional 150 neutral queries based on daily scenarios. We chose several widely used LLMs for evaluation, and tested them on the EQ+ dataset to assess the impact of adding EVs on their performance.
In the following experiments, unless specifically mentioned, we used the base emotion vector (EVbase) and applied different scalar factors to modulate the intensity. These variations were then applied to different models, and corresponding responses were generated for each query in EQ+ dataset. The full names of the models used in the following experiments are listed in Appendix A.
@@ 4.1 Sentence Fluency and Topic Adherence
Perplexity measures the fluency of a sentence based on a language model’s probability distribution over the next token. A lower perplexity indicates better fluency. To isolate the effects of applying EVs to hidden states under emotional conditioning, we used a separate pretrained model, Llama 3.1Dubey et al. (2024), to compute perplexity for each sentence, which is concatenated by the query and response. The final perplexity metrics are averaged on each sentence generated by the corresponding model. Details are shown in Appendix C.1
@@
Table 2: Perplexity scores for different models with $EV^{\text{base}}$ conditioning. $n*EV^{\text{base}}$ means that we apply n times of $EV^{\text{base}}$ to the model. When steering the $EV^{\text{base}}$ to the model shown as 1, we substitute $EV_{l}^{e_{k}}$ with $n*EV^{\text{base}}$.
Table 2 illustrates that the incorporation of emotional vectors (EV) has a negligible impact on sentence fluency across different models. While some models exhibit a slight decrease in fluency when EV is applied (e.g., Llama3.1 and Llama2 with 1EV), the magnitude of these decreases is minimal. Conversely, several models demonstrate an improvement in fluency under specific EV conditions, such as Llama3.1 with 2EV and baichuan2 with 2EV. These instances suggest that the addition of EV does not significantly compromise sentence fluency and can be effectively integrated into models.
Topic Adherence
For a chatbot, the consistency of answering questions is a very important indicator. The model’s answers should cover the same topics as the user’s questions. We call this ability "Topic Adherence". As modern models become more powerful, answers may not only cover user questions, but also have related extensions. Therefore, it is not appropriate to use traditional classification models for evaluation. Therefore, we choose to use GPT-4o-mini for evaluation. The specific evaluation prompts are given in the appendix C.2.
As shown in Table 3, most models retain very high topic adherence (almost the same as the topic adherence of the original answer) after EV is applied to the model. Models such as llama2, Qwen2.5 demonstrates very high robustness. llama3.1’s topic adherence decreases when applying EV because of the effectness when extracting the EV.
@@
Table 3: Topic Adherence scores for different models with $EV^{\text{base}}$ conditioning.

### v2

@@ Abstract
Purpose: Emotion is a fundamental component of human communication, shaping understanding, trust, and engagement across domains such as education, healthcare, and mental health. While large language models (LLMs) exhibit strong reasoning and knowledge generation capabilities, they still struggle to express emotions in a consistent, controllable, and contextually appropriate manner. This limitation restricts their potential for authentic human–AI interaction.
Methods: We propose a controllable emotion generation framework based on Emotion Vectors (EVs)—latent representations derived from internal activation shifts between neutral and emotion-conditioned responses. By injecting these vectors into the hidden states of pretrained LLMs during inference, our method enables fine-grained, continuous modulation of emotional tone without any additional training or architectural modification. We further provide theoretical analysis proving that EV steering enhances emotional expressivity while maintaining semantic fidelity and linguistic fluency.
Results: Extensive experiments across multiple LLM families show that the proposed approach achieves consistent emotional alignment, stable topic adherence, and controllable affect intensity. Compared with existing prompt-based and fine-tuning-based baselines, our method demonstrates superior flexibility and generalizability.
Conclusion: Emotion Vector (EV) steering provides an efficient and interpretable means of bridging rational reasoning and affective understanding in large language models, offering a promising direction for building emotionally resonant AI systems capable of more natural human–machine interaction.
keywords:
Latent Representation Fusion, Emotion Generation, Emotion Control, Vector-based Representation, Controllable Text Generation
@@ 1 Introduction
Emotion elevates us from mere information processors to beings with preferences, attachments, empathy, and the capacity to create meaning. It serves not merely as a tool, but as the fundamental means through which we experience life, comprehend the world, and establish profound connections with it. For instance, Education, healthcare, and mental health are among the most socially consequential domains of human life, where affective communication with various emotions is essential not only for solving practical problems but also for ensuring human well-being. In these domains, the role of emotion is deeply embedded. In education, learning outcomes are influenced not simply by the transmission of knowledge but by the emotional environment that sustains engagement and curiosity; a teacher’s encouragement or patience can significantly shape students’ motivation and persistence , which is show the effection of positive emotions [1, 2]. In healthcare, extensive research has demonstrated that physicians’ emotional engagement and empathic communication improve patient adherence, satisfaction, and even clinical recovery trajectories [3, 4]. Likewise, in mental health and companionship settings, the capacity for emotional attunement is not an accessory but a prerequisite for meaningful support [5, 6]. A counseling exchange devoid of empathy may satisfy informational needs yet leave the user’s deeper concerns unresolved. These examples underscore a critical principle: emotion is not ancillary to human-centered interaction but a central determinant of its effectiveness.
At the same time, the demand for such emotionally enriched interactions has been steadily growing with the development of modern society [7, 8]. Yet Relying solely on humans to provide stable, high-quality affective support is becoming increasingly untenable: human providers are inherently variable in availability and emotional consistency [9, 10, 11], and the cost of delivering continuous, personalized care at scale is prohibitively high [12], while these consistent emotional support is essential [13, 14]. This challenge between rising expectations for emotional value and the practical limits of human resources calls for technological solutions that can complement and extend human capacity.
The advent of large language models (LLMs) has opened new opportunities for intelligent systems to enter these domains at scale. Applications range from AI tutors that adaptively guide students [15], to clinical decision-support systems that assist physicians [16, 17, 18], to conversational agents offering mental health interventions or companionship [19, 20]. Early deployments have yielded promising results: LLM-powered tutoring has been shown to foster individualized learning experiences [21]; clinical support tools can streamline patient communication and diagnostic reasoning [22]; and dialogue agents for psychological support have increased accessibility to low-cost interventions [23]. Such progress demonstrates the transformative potential of LLMs in addressing some of society’s most pressing challenges.
Yet, despite their successes, current LLMs are fundamentally limited in their ability to engage with users on an emotional level. Most models generate content that is either affect-neutral, inconsistent in tone, or uncontrolled in affective orientation [24]. This shortcoming has been documented across multiple application studies: educational chatbots often fail to sustain motivational discourse [25]; medical assistants provide clinically accurate but emotionally detached responses; and mental health chatbots, while helpful in offering structured advice, lack the warmth, empathy, or reassurance characteristic of human counselors [26]. Moreover, existing reseaches have pointed out the concerns on the AI’s empathy, capabilities, safety, and human involvement in mental healthcare [27].
Such deficiencies not only reduce user trust but also constrain the overall efficacy of these systems [28]. Indeed, a purely factual but emotionally sterile response often leaves users dissatisfied, much as a technically competent but emotionally indifferent teacher, doctor, or counselor would fail to meet human expectations [29].
The consequences of this limitation are nontrivial. Consider the classroom, where an AI tutor might successfully solve a mathematical problem but fail to offer encouragement to a struggling student, thereby missing an opportunity to build resilience and confidence. In clinical care, an AI assistant may inform a patient of treatment side effects accurately yet omit the reassurance that a physician might naturally provide to alleviate anxiety. In psychological support, a chatbot may suggest coping strategies but without the empathetic validation that reassures individuals of their worth and shared humanity. In each of these cases, the absence of emotional intelligence constrains the system’s ability to produce meaningful, lasting  [30].
Aiming to evolve LLMs from problem-solving instruments into genuine human-centered collaborators, the capacity for controllable, consistent, and contextually appropriate emotional expression must be developed. Unlike the stochastic or incidental emotional cues sometimes present in current LLM outputs [31, 32], such emotion expression must be deliberate and adaptive, aligning with users’ affective states and situational demands. Crucially, this ability must not be hard-coded or manually scripted but rather systematically controllable within the generative process, enabling flexible, reliable, and self-consistent affective behavior. Achieving this capability is key to unlocking deeper integration of LLMs in education, healthcare, and mental health—domains where emotional attunement is as critical as cognitive competence.
@@
To overcome these challenges, we propose an elegant but effective method for the controllable emotional and affective expressions LLMs. Our approach offers a universal solution that allows fine-grained control over the emotional tone and sentiment of generated text, without compromising its fluency or coherence. It only needs to extract the ”Emotion Vector” used by the LLM to express basic emotions with simple prompts. This EV is then applied during inference to guide and control the emotional qualities of the generated output. Comprehensive evaluations across a range of LLM architectures confirm the consistency and stability of the resulting emotional expressions. This demonstrated universality addresses a key limitation of previous approaches, which were often constrained to specific models or datasets. Ultimately, this work provides valuable insights into how to equip LLMs with a reliable capacity for generating contextually appropriate emotional responses with minimal computational overhead.
2 Related works
Emotional Representations and Dialouge Systems
To create agents or dialogue systems that simulate human expression, a significant body of research has focused on understanding and representing emotions as fundamental aspects of human communication [33, 34]. Various theoretical frameworks and computational models have been explored to capture the multifaceted nature of emotions, including categorical approaches [35, 36] (e.g., discrete emotions like joy, sadness, anger [37]) and dimensional approaches [38] (e.g., valence-arousal scales [39]). These representations serve as the building blocks for infusing artificial systems with the ability to perceive, process, and generate emotional content, aiming to enable more natural and empathetic interactions with users. WhileZhou et al. [40] and Song et al. [41] proposed a way of Emotion Embedding to make the model ”has” the emotion, where, models were forced to install a module to generate emotions. However, most methods are too complex or requires further training. To achieve an effective emotion system, it is essential for the model to have precise, quantifiable control over emotions, as well as a flexible, plug-and-play module that can be seamlessly integrated as needed. It should also be consistent along the whole dialog.
Instruct tuning and prompt based emotional control
@@ 2 Related works
A significant body of work has focused on leveraging fine-tuning or prompt techniques for LLMs. Chen et al. [42], Chen et al. [43] and Zheng et al. [44] explored fine-tuning approaches to cultivate empathetic behavior in LLMs for psychological counseling and emotional supports. However, althrough instruct-tuning models have relatively good performance, they are often inflexible and struggle to adapt to a wide range of applications and model architectures, due to their predefined emotion categories or fixed sets of emotional datasets[45, 46]. Moreover, prompting strategies have also been used to elicit emotions without model modification. Li et al. [47], Wang et al. [48], Li et al. [49]However, prompting depends on elaborate templates and external evaluation modules to maintain effectiveness.
Inference-Time Vectors Editing
Recent studies have explored editing the internal representations of language models to achieve controlled generationDekoninck et al. [50], Liu et al. [51], Li et al. [52]. They uses latent steering vectors that enable semantic or stylistic shifts by modifying hidden activations. However, while they can realize controlable generation, these methods mainly focuses on the last token position during extraction and lacks global significanceTodd et al. [53]. It is difficult to apply to tasks such as emotions that require high generalization. Most control vector-related work is sentence-level controlSubramani et al. [54], and requires training, focusing only on regulating the model’s output for a single sentence. There has not been much success in achieving global control, which is essential for tasks like emotion control. A good emotion control system should be global, as this is necessary for building an effective emotion system.
Our Position
In contrast to the above paradigms, our method extracts reusable and efficient Emotion Vectors (EVs) by comparing model responses to emotion-inducing and neutral prompts. It is fully unsupervised, highly robust and controllable, requiring no training or architecture changes and is global consistent. EVs provide continuous and fine-grained control over emotional intensity through scalar scaling, enabling broad applicability across model families. Compared to previous approaches, EV offers a more general and efficient mechanism for emotion modulation in LLMs.
@@ 3 Method
Figure 1: Overview of the Emotion Vector (EV) pipeline. The figure follows the two-stage workflow used in our paper. EV extraction: For each target emotion $e_{k}$, the model is run on EmotionQuery with an emotion-conditioned prompt and a neutral prompt. At every transformer block $i$, we compute the token-averaged hidden outputs and take their difference $\Delta O_{i}^{(e_{k})}$; averaging over NN queries yields a layer-wise vector $\mathrm{EV}_{i}^{(e_{k})}=\tfrac{1}{N}\sum_{n=1}^{N}\Delta O_{i,n}^{(e_{k})}$. EV steering at inference: During generation, we inject the EV into the residual stream of every attention block, modifying the hidden state as $\hat{H}_{i}=H_{i}+\alpha\,\mathrm{EV}_{i}^{(e_{k})}$ (or $\alpha\,\mathrm{EV}^{\text{base}}$), and propagate the modified activations through subsequent self-attention/FFN blocks for each token. The scalar $\alpha$ provides continuous control of emotional intensity, and EVs can be combined additively if needed. This plug-and-play procedure leaves all model parameters frozen, yet steers the network toward the desired emotional direction while preserving semantic content.
@@ 3.1 Constructing Emotion Vectors
To capture the emotional factors and semantics for LLM, a specialized dataset is designed and constructed to elicit specific emotional responses, referred to as EmotionQuery. The dataset consists of 500 queries, with 100 queries generated for each of five emotional states derived from the basic emotion models[55]: joy, anger, disgust, fear, and sadness to provoke the corresponding emotional reactions. The queries were generated by a GPT-4o-mini[56]. A more detailed description of the dataset and query construction process can be found in the Appendix B.1.
@@
$\Delta O_{l}^{e_{k}}=\bar{O}_{l}^{\text{emotion}(e_{k})}-\bar{O}_{l}^{\text{neutral}},$ (2)
@@ 3.2 Steering Emotion Vectors
$EV^{\text{base}}=\frac{1}{K}\sum_{k=1}^{K}EV^{e_{k}},$ (6)
@@
4 Theoretical Rationale
To complement the empirical evidence, we provide here a concise theoretical rationale for why injecting layerwise Emotion Vectors (EVs) into the hidden states of a Transformer language model can reliably control emotional expression while preserving semantic fidelity. A full and rigorous proof is given in Appendix C.
Setting.
Consider a pretrained Transformer with $L$ layers. Let $H_{l}\in\mathbb{R}^{d}$ denote the mean-pooled hidden representation at layer $l$, with the update
$H_{l+1}=f_{l}(H_{l}),\quad l=0,\dots,L-1.$ (7)
The model output logits are
$z=W_{o}H_{L}+b\in\mathbb{R}^{V},$ (8)
which define the next-token distribution via a softmax.
An Emotion Vector for emotion $e$, denoted $EV^{(e)}_{l}$, is constructed as the difference between the mean hidden states produced by emotion-inducing and neutral prompts with matched semantics:
$EV^{(e)}_{l}\equiv\mathbb{E}\big[\overline{O}^{(e)}_{l}-\overline{O}^{(\mathrm{neutral})}_{l}\big$]. (9)
At inference time, we inject a scaled perturbation $\alpha EV^{(e)}_{l}$ at each layer,
$\widehat{H}_{l}=H_{l}+\alpha EV^{(e)}_{l}$, (10)
with $\alpha\in\mathbb{R}$ controlling the strength of emotional modulation.
Readout Functionals.
To disentangle emotion and semantic effects, we define two linear readouts:
$g(z)$ $=w_{e}^{\top}z,$ $\text{(emotion readout score for target emotion $e$)},$
$s(z)$ $=u^{\top}z,$ $\text{(semantic/topic adherence readout)}.$
Here $w_{e}\in\mathbb{R}^{V}$ is a fixed direction associated with the classifier or lexical indicator of emotion $e$, while $u\in\mathbb{R}^{V}$ corresponds to a direction sensitive to semantic or topical consistency. This formalization allows us to give a definite conclusion that EV injection reliably increase $g(z)$ without significantly altering $s(z)$.
Key Findings.
A first–order Taylor expansion of the network mapping gives
$\Delta z\approx\alpha\sum_{l=0}^{L-1}J_{l}EV^{(e)}_{l}$, (11)
where $J_{l}$ is the Jacobian $\partial z/\partial H_{l}$. This simple relation leads to three main theoretical guarantees:
  1. 1.
    Monotonic Emotion Gain. If the Fisher-discriminant direction of each layer aligns on average with the EV, then $\mathbb{E}[\Delta g]\propto\alpha>0$, i.e., small positive $\alpha$ monotonically increases the target emotion score. This provides a principled explanation for the empirically observed rise in emotion probability and confidence under $1\times$ and $2\times$ scaling.
  2. 2.
    Semantic Preservation. Because EVs are constructed from pairs of prompts with identical semantics but different emotions, their projection onto the semantic gradient $u^{\top}J_{l}$ is approximately zero. Consequently
    $\mathbb{E}[u^{\top}\Delta z]\approx 0,$ (12)
    showing that topic adherence is maintained and perplexity remains nearly unchanged.
  3. 3.
    Linear Controllability and Additivity. The linear dependence on $\alpha$ implies that emotion intensity grows proportionally to the scaling factor and that multiple emotions can be combined additively, $\sum_{k}\alpha_{k}EV^{(e_{k})}_{l}$, with predictable effects.
Robustness.
Distributing small shifts across all layers yields a favorable signal-to-noise ratio: aligned signals accumulate as $O(L)$ while unaligned noise grows only as $O(\sqrt{L})$. This explains why full-layer EV injection keeps topic coherence high and fluency stable, and why very large $\alpha$ may eventually cause saturation, as observed in $4\times$ experiments.
A complete mathematical treatment, including precise conditions and proofs of all propositions, is deferred to Appendix C.
@@ 5 Experiments
Guided by the theoretical analysis in Section 4, we empirically evaluate three key questions derived from our framework: (i) Controllability — whether emotion vector (EV) steering reliably induces the intended emotional tone in generated outputs; (ii) Semantic preservation — whether the EV injection preserves the original semantics and fluency of the sentences; and (iii) Linear controllability — whether emotional intensity increases predictably with the scaling factor $\alpha$. These correspond directly to the theoretical properties established in Section 4, namely monotonic emotion gain, semantic stability, and linear additivity.
To examine these questions, we evaluate on the EmotionQuery+ (EQ+) dataset, which extends the original EmotionQuery dataset with additional neutral and emotion-conditioned prompts. Specifically, EQ+ contains 50 queries for each of the five basic emotions (joy, anger, disgust, fear, and sadness) and 150 additional neutral queries representing daily scenarios, totaling 400 queries in all. The construction details of EQ+ are provided in Appendix B.2.
Unless otherwise specified, we use the base emotion vector ($\mathrm{EV}^{\text{base}}$)—the mean of all emotion-specific vectors—and apply different scalar factors $\alpha$ to modulate emotional intensity. For emotion-specific analyses, $\mathrm{EV}^{\text{base}}$ is replaced with the corresponding $\mathrm{EV}^{(e)}$. We evaluate several representative large language models (see Appendix A for full model names) and generate responses for every query in the EQ+ dataset under each condition.
@@ 5.1 Sentence Fluency and Topic Adherence
Perplexity measures the fluency of a sentence based on a language model’s probability distribution over the next token. A lower perplexity indicates better fluency. To isolate the effects of applying EVs to hidden states under emotional conditioning, we used a separate pretrained model, Llama 3.1[57], to compute perplexity for each sentence, which is concatenated by the query and response. The final perplexity metrics are averaged on each sentence generated by the corresponding model. Details are shown in Appendix D.1.
Table 2(a) illustrates that the incorporation of emotional vectors (EV) has a negligible impact on sentence fluency across different models. While some models exhibit a slight decrease in fluency when EV is applied (e.g., Llama3.1 and Llama2 with 1EV), the magnitude of these decreases is minimal. Conversely, several models demonstrate an improvement in fluency under specific EV conditions, such as Llama3.1 with 2EV and baichuan2 with 2EV. These instances suggest that the addition of EV does not significantly compromise sentence fluency and can be effectively integrated into models.
Topic Adherence
For conversational agents, maintaining consistency between user queries and model responses is a crucial quality indicator. A model should generate answers that remain aligned with the user’s intended topic—a capability we refer to as Topic Adherence. As modern large language models grow increasingly capable, their responses often include not only direct answers but also relevant extensions or elaborations. This makes traditional classification-based evaluation methods inadequate for measuring topic consistency. To better capture this nuance, we employ GPT-4o-mini as an evaluator, using carefully designed prompts detailed in Appendix D.2. As shown in Table 2(b), most models maintain remarkably high topic adherence after applying the EV, with results comparable to those of the original, unmodified responses. Models such as Llama-2 and Qwen-2.5 exhibit particularly strong robustness under EV steering. In contrast, Llama-3.1 shows a slight degradation in topic adherence, which can be attributed to the relatively large norm of its extracted EV. This excessive magnitude perturbs the model’s later decoding process, leading to minor semantic deviations in the generated responses.
@@
((a)) Perplexity scores for different models with $EV^{\text{base}}$ conditioning. $n*EV^{\text{base}}$ means applying $n$ times of $EV^{\text{base}}$. When steering $EV^{\text{base}}$ as in Eq. 4, we substitute $EV_{l}^{e_{k}}$ with $n*EV^{\text{base}}$.
@@
((b)) Topic Adherence scores for different models with $EV^{\text{base}}$ conditioning.